Corrections to Probab. Theory Relat. Fields 100, 191-203 (1994) Martin Hildebrand

1. In the proof of Lemma 4, on p. 197, twice the expression (1 + 2/(ϵ/2)^{1/(k−1)}n^{(k−2)/(k−1)}))^{k−1} should be replaced with (1 + 2/((ϵ/2)^{1/(k−1)}n^{(k−2)/(k−1)})))^{k−1}

2. The display in the last sentence of the proof of Lemma 4 does not hold if k = 2, and the proof of the lemma is incomplete for k = 2. Lemma 4 is used to prove Theorem 1. However, Theorem 1 in the case k = 2 can be readily proven by Theorem 2 (which is proved without Theorem 1 or Lemma 4).

The author regrets these errors.